OpenStudy (anonymous):

e^x-4=e^-x (Solve) I came up with 0

OpenStudy (anonymous):

u mean e^(x-4) = e^(-x) ?.

OpenStudy (anonymous):

x=-2

OpenStudy (anonymous):

I don't get 0 as the answer. \[e^{x-4}=e^{-x}\]\[\frac{e^{x-4}}{e^{-x}}=1\]\[e^{2x-4}=e^{0}\]\[2x-4=0\]\[2x=4\]\[x=2\]

OpenStudy (anonymous):

How have you gone about solving the question koons4?

OpenStudy (anonymous):

I must of not finished it out, because I only got 2x-4=0 That is where I think I messed up.

OpenStudy (anonymous):

No worries, good luck :)

OpenStudy (anonymous):

Trust me I need it!